# Implicit Bias and Invariance: How Hopfield Networks Efficiently Learn Graph Orbits

### v1

@@ 1 Introduction
Here, we analyze the emergence of invariance arising implicitly during training in Hopfield networks (HNs) (hopfield1982neural), which represent arguably the simplest example of an Associative Memory. Building on classical ideas  rosenblatt1958perceptron; willshaw1969non; amari1972learning; little1974existence; pastur1977exactly, HNs are recurrent neural networks consisting of $n$ linear-threshold McCulloch–Pitts neurons mcculloch1943logical that can store binary patterns as distributed memories in the form of fixed-point attractors of its recurrent dynamics. In the literature, HNs are usually associated with a particular Hebbian learning scheme called the “Outer-Product Rule", but for the purposes of this work we also consider other standard training methods. This setting is intentionally minimal so that we can focus on developing novel mathematical tools for understanding generalization in a classical architecture. As data symmetry is not made explicit in this model, any invariance must arise from the interplay of the group structure in the data and the implicit bias of the learning rule in question. More specifically, and inspired by hillar2018robust; hillar2021hidden, this paper studies whether or not standard learning rules and objectives, notably minimization of the energy flow (MEF) mpf-hopf, a tractable convex loss, can learn the isomorphism class of a graph from a small, random subset. Our key findings are as follows.
@@
    Polynomial sample complexity suffices for orbit generalization. Suppose $D\subset\{0,1\}^{n}$ is strictly memorizable with min-norm parameter $\mbox{$\theta$}^{*}$, $\|{\bm{x}}\|_{0}\leq m$ for all ${\bm{x}}\in D$, and let $\mathcal{D}$ be a distribution supported on $D$. We prove $N=\tilde{\Omega}\!\left(n\|\mbox{$\theta$}^{*}\|^{2}m\epsilon^{-2}\right)$ random samples suffices for both HSVM and MEF to memorize new samples with probability at least $1-\epsilon$ (Theorem 3.2). These theoretical results corroborate the empirical “few-shot-to-orbit” phenomenon we observe (Figs. 1, 2, 5, 7, 8, 9). Moreover, specializing to isomorphism classes this result implies a polynomial sample complexity in the number of vertices $v$, supporting a conjecture in hillar2021hidden for the case of cliques. In this case, the critical number of samples for generalization appears to coincide with when these networks solve the Hidden Clique Problem (HCP) dekel2014finding; see Fig. 5 in Appendix C.2. These critical sample counts can also be used to differentiate between combinatorial classes of graphs (see Fig. 2).
@@ 1.1 Related work
Capacity of Hopfield networks. The capacity of a Hopfield network depends on the learning rule and the structure of the data. For dense, uncorrelated random patterns under Hebbian learning, the statistical–mechanics analysis of (amit1985storing) gives the classic linear law: reliable retrieval up to approximately $0.138$ patterns per neuron with subsequent refinements via replica methods (gardner1988space; krauth1989storage). Coding-theoretic analyses further show that, for Hebbian constructions and exact recovery of randomly chosen patterns, one typically cannot exceed $n/(4\ln n)$ (MR901677) memories. More generally, Cover’s classical bound cover1965geometrical restricts the capacity for exact storage of dense, random data to only $2n$. Nonetheless, superlinear capacity is achievable for certain structured datasets. For example, sparse data having few active neurons can yield an increase of capacity to nearly quadratic in $n$ (tsodyks1988enhanced; amari1989characteristics). Additionally, robust exponential memory has been observed for particular examples of group structured data (hillar2018robust; hillar2021hidden); in particular, for storing all $k$-clique graphs and their hypergraph analogues. Our work builds on the observations of (hillar2018robust; hillar2021hidden) by proving that all graph isomorphism classes are memorizable.
@@
Modern Hopfield Networks. A line of recent investigation has sought to increase the capacity and retrieval properties of Hopfield networks by changing the energy function. Dense Associative Memories (DAMs) replace the classical quadratic energy with higher-order polynomial interactions, resulting in a capacity that scales polynomially with neuron count (krotov2016dense; horn1988capacities). Building on this, Modern Hopfield Networks (MHNs) introduced a log-sum-exp energy function that allows the capacity to grow exponentially (ramsauer2020hopfield; demircigil2017model). Our work provides a complementary perspective to these advancements by showing classical HNs can achieve exponential capacity by capturing the symmetry of the data in its parameters. Preliminary experiments suggest that DAMs do not generalize well in this setting (see Fig. 6, Appendix C.3).
@@
Generalization beyond the training set for HNs. Theoretical study of classical HNs primarily focuses on storage limits, basins of attraction, and noise robustness around memorized patterns, rather than sample–complexity guarantees for generalization to patterns outside the training set. Earlier analyses of concept generalization in classical HNs investigate when networks capture latent data regularities (fontanari1990generalization). More recently, Random-Features Hopfield Models (RFHMs) exhibit learning phase transitions and even retrieval of previously unseen examples (negri2023rfhm; kalaj2024rfhmgen). These results complement but are distinct from our own; in particular, they do not provide sample–complexity bounds or analyze the emergence of invariance induced by a symmetry present in the data. In addition, while these results primarily use techniques from statistical physics, here we leverage tools from statistical learning theory.
@@
Emergent invariance through data augmentation and feature averaging. One perspective on data augmentation is as orbit averaging over a symmetry group; in particular, empirical risk minimization on augmented data is equivalent to averaging features or predictions along group orbits. This has been shown to induce approximate invariance and reduces estimator variance (chen2020gtda). From a kernel perspective, augmentation decomposes into first-order invariant feature averaging plus a second-order variance regularizer (dao2019kernelaugment). Enforcing invariance through this averaging method yields provable generalization benefits in the context of invariant kernel regression (elesedy2021provably). Beyond static estimators, recent results show that with full group augmentation deep ensembles become equivariant in expectation at all training times in the infinite-width limit (gerken2024emergent) and that the expected predictions of group-convolutional networks match those of data-augmented conventional networks throughout training (eNTK_arxiv2024). While these results assume explicit invariance, either through architectural design or by averaging over the full group orbit, here we ask whether simple learning rules can implicitly recover approximate invariance from small random sample of elements.
@@
Implicit bias. A large body of work shows that, even without explicit regularization, certain learning dynamics have a preference for particular solutions. In particular, for classification using the logistic loss, gradient descent drives the parameter norm to infinity while the parameter direction converges to the max–margin classifier (soudry2018the; ji2019risk; nacson2019convergence). Our work leverages these results in order to show that standard learning rules for HNs are implicitly biased towards learning invariant representations when trained on group data.
@@ 2 Preliminaries
we consider a Hopfield network hopfield1982neural with asynchronous dynamics but do not restrict ourselves to Hebbian learning. To this end, let $\textnormal{Sym}_{0}^{n}\subset\mathbb{R}^{n\times n}$ denote the set of symmetric, real, $n\times n$ matrices whose diagonal entries are zero, and let $\Theta=\textnormal{Sym}_{0}^{n}\times\mathbb{R}^{n}$. Clearly $\Theta$ is a convex vector subspace. We introduce the energy function $E:\{0,1\}^{n}\times{\bm{\Theta}}\rightarrow\mathbb{R}$ defined as
@@
for all $t\geq 1$ and $j\in[n]$. For any input ${\bm{x}}\in\{0,1\}^{n}$, this sequence converges in finite time to a fixed point bruck-hopfiel-convergence. We define the input-output map of the Hopfield network, denoted $H_{\theta}:\{0,1\}^{n}\rightarrow\{0,1\}^{n}$ as follows: given an input $\bm{x}$}, the output $H_{\theta}({\bm{x}})$ is the attractor or fixed point of 2 reached when initialized with ${\bm{x}}(0)={\bm{x}}$. If $H_{\theta}({\bm{x}})={\bm{x}}$, then $\bm{x}$} is a fixed point of the recurrent dynamics, and furthermore, we say that $H_{\theta}$ has memorized $\bm{x}$}. A sufficient condition for $H_{\theta}$ to memorize $\bm{x}$} is $E({\bm{x}};\mbox{$\theta$})<E({\bm{x}}^{\prime};\mbox{$\theta$})$ for all ${\bm{x}}^{\prime}\in\mathcal{N}({\bm{x}})$, where here $\mathcal{N}({\bm{x}})$ denotes the set of all binary vectors a Hamming distance of exactly one from $\bm{x}$}. If $\theta$ satisfies this property, we say that $H_{\theta}$ strictly memorizes $\bm{x}$}. We also denote the action of a permutation matrix ${\bm{P}}\in\mathcal{P}_{n}$ on the parameters of a Hopfield network as ${\bm{P}}\mbox{$\theta$}\coloneq({\bm{P}}^{T}{\bm{W}}{\bm{P}},{\bm{P}}^{T}{\bm{b}})$.
@@
let $\mathcal{S}\subset\{0,1\}^{n}$, we say that $H_{\theta}$ memorizes $\mathcal{S}$ if it memorizes all ${\bm{x}}\in\mathcal{S}$. There are many methods Hertz91; Tolmachev20 that have been proposed to train networks to memorize a set $\mathcal{S}$ (Appendix A.2). We focus on minimization of the energy flow (MEF) mpf-hopf; hillar2018robust; hillar2021hidden and study its implicit bias. If ${\bm{x}}\in\{0,1\}^{n}$ and $j\in[n]$, let ${\bm{x}}^{(j)}\in\{0,1\}^{n}$ satisfy $x_{l}\neq x^{(j)}_{l}$ iff $l=j$. We define the energy flow loss as:
@@
For any given set of points $\mathcal{S}$, note that $L$ is nonnegative, infinitely differentiable, and is convex in $\Theta$. As a result, minimizing $L$ is a convex problem to which a wide variety of numerical techniques can be applied, including but not limited to variants of gradient descent (GD) as well as (approximate) second order methods such as L-BFGS liu1989limited. As long as $\mathcal{S}$ can be memorized, then sufficiently minimizing 3 will result in a network which memorizes $\mathcal{S}$. For further details on MEF learning for discrete recurrent neural networks, we refer the reader to hillar2021hidden and its inspiration from the theory of density estimation PhysRevLett.107.220601.
@@ 3 Implicit bias, minimum norm memorizers and generalization
Applying (soudry2018the, Thm.3) it can be proved that this sequence of GD iterates converges in direction to the solution of (7).
@@
((soudry2018the, Thm.3) adapted to our setting) Assume $\mathcal{S}$ can be strictly memorized, let ${\bm{\omega}}^{*}=\textnormal{HSVM}(\mathcal{S}_{N})$, ${\bm{\omega}}(0)\in\mathbb{R}^{q}$ be arbitrary and ${\bm{\omega}}(t)$ be generated for all $t\in\mathbb{N}_{\geq 1}$ as per (8). There exists a choice of step size $\eta$ such that ${\bm{\omega}}(t)={\bm{\omega}}^{*}\log(t)+\rho(t)$ for all $t\in\mathbb{N}_{\geq 1}$, where $\rho(t)$ grows as $\|\rho(t)\|=O(\log(\log(t)))$. Moreover, $\lim_{t\rightarrow\infty}\tfrac{{\bm{\omega}}(t)}{\|{\bm{\omega}}(t)\|}=\tfrac{{\bm{\omega}}^{*}}{\|{\bm{\omega}}^{*}\|}.$
@@
To prove Theorem 3.2 we combine a vector contraction inequality (vector-contrac, Corollary 1) with Rademacher bounds, see e.g., (FoML, Theorem 3.3), we refer the reader to Appendix B.1 for a full proof. It is worth emphasizing that Theorem 3.2 implies any dataset $D$ which can be strictly memorized, can be at least nearly strictly memorized using only a polynomial number of samples. In Section 4.3 we take preliminary steps towards relaxing this statement, from memorizing samples drawn from $\mathcal{D}$ with high probability, to memorizing the set $D$ itself with high probability. Finally, we remark that the MEF bound implies gradient descent may require exponentially many iterations to converge directionally to the max-margin solution. We hypothesize that this is a tail phenomenon: once the weights are approximately aligned, all points are classified with a significant margin and their loss contributions become exponentially small.
@@ 4.2 Experiments on graph data
To experimentally assess storage across isomorphism classes we study several classes of graphs; namely, cliques, bipartite, chain, cycle, Johnson, and circulant. Clique graphs, or more specifically $k$-cliques, contain a fully connected subset of $k$ vertices while the remaining vkv-k vertices are isolated. Bipartite graphs split the $v$ vertices into two equally sized groups with all possible inter-group edges present and no intra-group edges. Paley graphs connect vertices $l$ and $j$ when (lj)(l-j) is a quadratic residue mod $v$, as per NetworkX networkx; cf. bollobás2001randomgraphs. Definitions of other graphs are readily found, e.g., on Wikipedia. We remark that extensive experiments for cliques are already provided in hillar2018robust; hillar2021hidden; we include them here again for comparison and completeness. We remark that we selected these classes due to the ease with which we are able to sample from them and emphasize that these families are representative rather than special. Indeed, we observe similar behavior for many other graph isomorphism classes, including random graphs. Code and experiments are available at https://github.com/hopnetorbit/HopfieldNetworksIsomorphism.
Fig. 1 shows test accuracy versus training sample size, with mean and min–max over 10 trials, for Hopfield networks trained by MEF, Perceptron, and Delta (the latter two used only as baselines; see Appendix A.2). For small graphs ($v=8$) we enumerate the full isomorphism class and report the true accuracy, i.e., the fraction of the class memorized. For larger graphs ($v=20$), accuracy is estimated on an independent random sample of 1000 graphs. We highlight two observations: (i) MEF appears to reach higher test accuracy with fewer samples relative to the other methods, despite all methods perfectly memorizing the training set. This suggests differing implicit biases or implicit bias strengths. (ii) For MEF and Delta, the sample size needed to memorize an isomorphism class is tiny relative to the class size, aligning with the findings in hillar2018robust; hillar2021hidden. Furthermore, the number of iterations was capped at 1000, suggesting that the exponential dependency in Theorem 3.2 is highly pessimistic. Finally, within our hyperparameter range, the Delta rule using Adam failed on the $k$-clique class, whereas MEF learned all classes and was insensitive to optimizer choice.

### v2

@@ 1 Introduction
Here, we analyze the emergence of invariance arising implicitly during training in Hopfield networks (HNs) (Hopfield, 1982), which represent arguably the simplest example of an Associative Memory. Building on classical ideas  Rosenblatt (1958); Willshaw et al. (1969); Amari (1972); Little (1974); Pastur and Figotin (1977), HNs are recurrent neural networks consisting of $n$ linear-threshold McCulloch–Pitts neurons McCulloch and Pitts (1943) that can store binary patterns as distributed memories in the form of fixed-point attractors of its recurrent dynamics. In the literature, HNs are usually associated with a particular Hebbian learning scheme called the “Outer-Product Rule", but for the purposes of this work we also consider other standard training methods. This setting is intentionally minimal so that we can focus on developing novel mathematical tools for understanding generalization in a classical architecture. As data symmetry is not made explicit in this model, any invariance must arise from the interplay of the group structure in the data and the implicit bias of the learning rule in question. More specifically, and inspired by Hillar and Tran (2018); Hillar et al. (2021), this paper studies whether or not standard learning rules and objectives, notably minimization of the energy flow (MEF) Hillar et al. (2012), a tractable convex loss, can learn the isomorphism class of a graph from a small, random subset. Our key findings are as follows.
@@
    Polynomial sample complexity suffices for orbit generalization. Suppose $D\subset\{0,1\}^{n}$ is strictly memorizable with min-norm parameter $\mbox{$\theta$}^{*}$, $\|{\bm{x}}\|_{0}\leq m$ for all ${\bm{x}}\in D$, and let $\mathcal{D}$ be a distribution supported on $D$. We prove $N=\tilde{\Omega}\!\left(n\|\mbox{$\theta$}^{*}\|^{2}m\epsilon^{-2}\right)$ random samples suffices for both HSVM and MEF to memorize new samples with probability at least $1-\epsilon$ (Theorem 3.2). These theoretical results corroborate the empirical “few-shot-to-orbit” phenomenon we observe (Figs. 1, 2, 5, 7, 8, 9). Moreover, specializing to isomorphism classes this result implies a polynomial sample complexity in the number of vertices $v$, supporting a conjecture in Hillar et al. (2021) for the case of cliques. In this case, the critical number of samples for generalization appears to coincide with when these networks solve the Hidden Clique Problem (HCP) Dekel et al. (2014); see Fig. 5 in Appendix C.2. These critical sample counts can also be used to differentiate between combinatorial classes of graphs (see Fig. 2).
@@ 1.1 Related work
Capacity of Hopfield networks. The capacity of a Hopfield network depends on the learning rule and the structure of the data. For dense, uncorrelated random patterns under Hebbian learning, the statistical–mechanics analysis of (Amit et al., 1985) gives the classic linear law: reliable retrieval up to approximately $0.138$ patterns per neuron with subsequent refinements via replica methods (Gardner, 1988; Krauth and Mézard, 1989). Coding-theoretic analyses further show that, for Hebbian constructions and exact recovery of randomly chosen patterns, one typically cannot exceed $n/(4\ln n)$ (McEliece et al., 1987) memories. More generally, Cover’s classical bound Cover (1965) restricts the capacity for exact storage of dense, random data to only $2n$. Nonetheless, superlinear capacity is achievable for certain structured datasets. For example, sparse data having few active neurons can yield an increase of capacity to nearly quadratic in $n$ (Tsodyks and Feigel’man, 1988; Amari, 1989). Additionally, robust exponential memory has been observed for particular examples of group structured data (Hillar and Tran, 2018; Hillar et al., 2021); in particular, for storing all $k$-clique graphs and their hypergraph analogues. Our work builds on the observations of (Hillar and Tran, 2018; Hillar et al., 2021) by proving that all graph isomorphism classes are memorizable.
@@
Modern Hopfield Networks. A line of recent investigation has sought to increase the capacity and retrieval properties of Hopfield networks by changing the energy function. Dense Associative Memories (DAMs) replace the classical quadratic energy with higher-order polynomial interactions, resulting in a capacity that scales polynomially with neuron count (Krotov and Hopfield, 2016; Horn and Usher, 1988). Building on this, Modern Hopfield Networks (MHNs) introduced a log-sum-exp energy function that allows the capacity to grow exponentially (Ramsauer et al., 2020; Demircigil et al., 2017). Our work provides a complementary perspective to these advancements by showing classical HNs can achieve exponential capacity by capturing the symmetry of the data in its parameters. Preliminary experiments suggest that DAMs do not generalize well in this setting (see Fig. 6, Appendix C.3).
@@
Generalization beyond the training set for HNs. Theoretical study of classical HNs primarily focuses on storage limits, basins of attraction, and noise robustness around memorized patterns, rather than sample–complexity guarantees for generalization to patterns outside the training set. Earlier analyses of concept generalization in classical HNs investigate when networks capture latent data regularities (Fontanari, 1990). More recently, Random-Features Hopfield Models (RFHMs) exhibit learning phase transitions and even retrieval of previously unseen examples (Negri et al., 2023; Kalaj et al., 2024). These results complement but are distinct from our own; in particular, they do not provide sample–complexity bounds or analyze the emergence of invariance induced by a symmetry present in the data. In addition, while these results primarily use techniques from statistical physics, here we leverage tools from statistical learning theory.
@@
Emergent invariance through data augmentation and feature averaging. One perspective on data augmentation is as orbit averaging over a symmetry group; in particular, empirical risk minimization on augmented data is equivalent to averaging features or predictions along group orbits. This has been shown to induce approximate invariance and reduces estimator variance (Chen et al., 2020). From a kernel perspective, augmentation decomposes into first-order invariant feature averaging plus a second-order variance regularizer (Dao et al., 2019). Enforcing invariance through this averaging method yields provable generalization benefits in the context of invariant kernel regression (Elesedy and Zaidi, 2021). Beyond static estimators, recent results show that with full group augmentation deep ensembles become equivariant in expectation at all training times in the infinite-width limit (Gerken and Kessel, 2024) and that the expected predictions of group-convolutional networks match those of data-augmented conventional networks throughout training (Marthaler et al., 2024). While these results assume explicit invariance, either through architectural design or by averaging over the full group orbit, here we ask whether simple learning rules can implicitly recover approximate invariance from small random sample of elements.
@@
Implicit bias. A large body of work shows that, even without explicit regularization, certain learning dynamics have a preference for particular solutions. In particular, for classification using the logistic loss, gradient descent drives the parameter norm to infinity while the parameter direction converges to the max–margin classifier (Soudry et al., 2018; Ji and Telgarsky, 2018; Nacson et al., 2019). Our work leverages these results in order to show that standard learning rules for HNs are implicitly biased towards learning invariant representations when trained on group data.
@@ 2 Preliminaries
we consider a Hopfield network Hopfield (1982) with asynchronous dynamics but do not restrict ourselves to Hebbian learning. To this end, let $\textnormal{Sym}_{0}^{n}\subset\mathbb{R}^{n\times n}$ denote the set of symmetric, real, $n\times n$ matrices whose diagonal entries are zero, and let $\Theta=\textnormal{Sym}_{0}^{n}\times\mathbb{R}^{n}$. Clearly $\Theta$ is a convex vector subspace. We introduce the energy function $E:\{0,1\}^{n}\times{\bm{\Theta}}\rightarrow\mathbb{R}$ defined as
@@
for all $t\geq 1$ and $j\in[n]$. For any input ${\bm{x}}\in\{0,1\}^{n}$, this sequence converges in finite time to a fixed point Bruck (1990). We define the input-output map of the Hopfield network, denoted $H_{\theta}:\{0,1\}^{n}\rightarrow\{0,1\}^{n}$ as follows: given an input $\bm{x}$}, the output $H_{\theta}({\bm{x}})$ is the attractor or fixed point of 2 reached when initialized with ${\bm{x}}(0)={\bm{x}}$. If $H_{\theta}({\bm{x}})={\bm{x}}$, then $\bm{x}$} is a fixed point of the recurrent dynamics, and furthermore, we say that $H_{\theta}$ has memorized $\bm{x}$}. A sufficient condition for $H_{\theta}$ to memorize $\bm{x}$} is $E({\bm{x}};\mbox{$\theta$})<E({\bm{x}}^{\prime};\mbox{$\theta$})$ for all ${\bm{x}}^{\prime}\in\mathcal{N}({\bm{x}})$, where here $\mathcal{N}({\bm{x}})$ denotes the set of all binary vectors a Hamming distance of exactly one from $\bm{x}$}. If $\theta$ satisfies this property, we say that $H_{\theta}$ strictly memorizes $\bm{x}$}. We also denote the action of a permutation matrix ${\bm{P}}\in\mathcal{P}_{n}$ on the parameters of a Hopfield network as ${\bm{P}}\mbox{$\theta$}\coloneq({\bm{P}}^{T}{\bm{W}}{\bm{P}},{\bm{P}}^{T}{\bm{b}})$.
@@
let $\mathcal{S}\subset\{0,1\}^{n}$, we say that $H_{\theta}$ memorizes $\mathcal{S}$ if it memorizes all ${\bm{x}}\in\mathcal{S}$. There are many methods Hertz et al. (1991); Tolmachev and Manton (2020) that have been proposed to train networks to memorize a set $\mathcal{S}$ (Appendix A.2). We focus on minimization of the energy flow (MEF) Hillar et al. (2012); Hillar and Tran (2018); Hillar et al. (2021) and study its implicit bias. If ${\bm{x}}\in\{0,1\}^{n}$ and $j\in[n]$, let ${\bm{x}}^{(j)}\in\{0,1\}^{n}$ satisfy $x_{l}\neq x^{(j)}_{l}$ iff $l=j$. We define the energy flow loss as:
@@
For any given set of points $\mathcal{S}$, note that $L$ is nonnegative, infinitely differentiable, and is convex in $\Theta$. As a result, minimizing $L$ is a convex problem to which a wide variety of numerical techniques can be applied, including but not limited to variants of gradient descent (GD) as well as (approximate) second order methods such as L-BFGS Liu and Nocedal (1989). As long as $\mathcal{S}$ can be memorized, then sufficiently minimizing 3 will result in a network which memorizes $\mathcal{S}$. For further details on MEF learning for discrete recurrent neural networks, we refer the reader to Hillar et al. (2021) and its inspiration from the theory of density estimation Sohl-Dickstein et al. (2011).
@@ 3 Implicit bias, minimum norm memorizers and generalization
Applying (Soudry et al., 2018, Thm.3) it can be proved that this sequence of GD iterates converges in direction to the solution of (7).
@@
((Soudry et al., 2018, Thm.3) adapted to our setting) Assume $\mathcal{S}$ can be strictly memorized, let ${\bm{\omega}}^{*}=\textnormal{HSVM}(\mathcal{S}_{N})$, ${\bm{\omega}}(0)\in\mathbb{R}^{q}$ be arbitrary and ${\bm{\omega}}(t)$ be generated for all $t\in\mathbb{N}_{\geq 1}$ as per (8). There exists a choice of step size $\eta$ such that ${\bm{\omega}}(t)={\bm{\omega}}^{*}\log(t)+\rho(t)$ for all $t\in\mathbb{N}_{\geq 1}$, where $\rho(t)$ grows as $\|\rho(t)\|=O(\log(\log(t)))$. Moreover, $\lim_{t\rightarrow\infty}\tfrac{{\bm{\omega}}(t)}{\|{\bm{\omega}}(t)\|}=\tfrac{{\bm{\omega}}^{*}}{\|{\bm{\omega}}^{*}\|}.$
@@
To prove Theorem 3.2 we combine a vector contraction inequality (Maurer, 2016, Corollary 1) with Rademacher bounds, see e.g., (Mohri et al., 2018, Theorem 3.3), we refer the reader to Appendix B.1 for a full proof. It is worth emphasizing that Theorem 3.2 implies any dataset $D$ which can be strictly memorized, can be at least nearly strictly memorized using only a polynomial number of samples. In Section 4.3 we take preliminary steps towards relaxing this statement, from memorizing samples drawn from $\mathcal{D}$ with high probability, to memorizing the set $D$ itself with high probability. Finally, we remark that the MEF bound implies gradient descent may require exponentially many iterations to converge directionally to the max-margin solution. We hypothesize that this is a tail phenomenon: once the weights are approximately aligned, all points are classified with a significant margin and their loss contributions become exponentially small.
@@ 4.2 Experiments on graph data
To experimentally assess storage across isomorphism classes we study several classes of graphs; namely, cliques, bipartite, chain, cycle, Johnson, and circulant. Clique graphs, or more specifically $k$-cliques, contain a fully connected subset of $k$ vertices while the remaining vkv-k vertices are isolated. Bipartite graphs split the $v$ vertices into two equally sized groups with all possible inter-group edges present and no intra-group edges. Paley graphs connect vertices $l$ and $j$ when (lj)(l-j) is a quadratic residue mod $v$, as per NetworkX Developers ; cf. bollobás2001randomgraphs. Definitions of other graphs are readily found, e.g., on Wikipedia. We remark that extensive experiments for cliques are already provided in Hillar and Tran (2018); Hillar et al. (2021); we include them here again for comparison and completeness. We remark that we selected these classes due to the ease with which we are able to sample from them and emphasize that these families are representative rather than special. Indeed, we observe similar behavior for many other graph isomorphism classes, including random graphs. Code and experiments are available at https://github.com/hopnetorbit/HopfieldNetworksIsomorphism.
Fig. 1 shows test accuracy versus training sample size, with mean and min–max over 10 trials, for Hopfield networks trained by MEF, Perceptron, and Delta (the latter two used only as baselines; see Appendix A.2). For small graphs ($v=8$) we enumerate the full isomorphism class and report the true accuracy, i.e., the fraction of the class memorized. For larger graphs ($v=20$), accuracy is estimated on an independent random sample of 1000 graphs. We highlight two observations: (i) MEF appears to reach higher test accuracy with fewer samples relative to the other methods, despite all methods perfectly memorizing the training set. This suggests differing implicit biases or implicit bias strengths. (ii) For MEF and Delta, the sample size needed to memorize an isomorphism class is tiny relative to the class size, aligning with the findings in Hillar and Tran (2018); Hillar et al. (2021). Furthermore, the number of iterations was capped at 1000, suggesting that the exponential dependency in Theorem 3.2 is highly pessimistic. Finally, within our hyperparameter range, the Delta rule using Adam failed on the $k$-clique class, whereas MEF learned all classes and was insensitive to optimizer choice.